OpenStudy (anonymous):

square root of x^2-x-4 =x+2 i got stuck on this one can anyone help me get through it with steps?

OpenStudy (anonymous):

\[\sqrt{x^2-x-4}=x+2???

OpenStudy (anonymous):

square both sides and solve

OpenStudy (anonymous):

i did i got stuck at x^2-x-4=x^2+4x+4

OpenStudy (anonymous):

combine like terms to one side of equation, notice x^2 will cancel out

OpenStudy (anonymous):

yes ill get 5x+8

OpenStudy (anonymous):

0=5x+8

OpenStudy (anonymous):

is the answer just -1.6?

OpenStudy (anonymous):

-8 = 5x -8/5 = x

OpenStudy (anonymous):

-1.6?